To construct a solenoid, you wrap insulated wire uniformly around a plastic tube 14cm in diameter and 50cm in length. You would like a 2.0❝A current to produce a 2.4❝kG magnetic field inside your solenoid.
Answer #1

apply magnitic fiedl due to solonoid as B = uo n i

where n = N/L

so B = uoNi/L

N = BL/uoi

N = 2400e-4 * 0.5/(4pie-7* 2)

N = 47770.7 turns

length=2pirN

L = 2*pi*0.07*47770.7

L = 21 Km
Answer #2

r=0,07m
h=0,50 m
I=2 A
B=2 400

B=mu0*N*I/h

mu0 --4pi*10powered-7
N -- number of turns on the solenoid
I -- current
h -- height of the solenoid
From here you get how many turns of wire you have
N=B*h/(mu0*I)

The length of the wire is I hope you agree
L=N*2*pi*r